ABC Ltd is evaluating to invest in two projects. Project X may yield a return of £1.5m with a probability of 20%, or a return of £5m with a probability of 60%. Project Y, instead, may earn a negative return of £3m with a probability of 70% or a positive return of £7m with a probability of 30%. 

How much is the expected return for project X and Y? [8 marks] 2) What is the standard deviation? [6 marks] 3) Compare and briefly discuss the expected return and risk of the projects.


1
Expert's answer
2021-07-08T14:10:02-0400

1.How much is the expected return for project X and Y?

Expected return = "\sum" (Returni"\times Probability"i)

For project X: (20% "\times" 1.5m) + (60% "\times" 5m)

= 300,000 + 3m

Expected return for project X is €3.3million

For project Y: (70% "\times" -3m) + (30% "\times" 7m)

= -2.1m + 2.1m

Expected return for project Y is €0


2) What is the standard deviation

"\sigma" = "\sqrt\frac{\sum(x-\mu)^{2}}{n-1}"

Project X "\sigma" = "\sqrt{(1.5m-3.3m)^{2}\times0.2 + (5m-3.3m)^{2}\times0.6}"

Project X "\sigma" = 1,543,372.93

Project Y "\sigma" = "\sqrt{(-3m-0)^{2}\times0.7 + (7m-0)^{2}\times0.3}"

Project Y "\sigma" = 4,582,575.70

3) Compare and briefly discuss the expected return and risk of the projects.

Project X has an expected return of €3.3m while project Y has an expected return of €0 on investment. Going by expected return, Project X poses has a higher expected return for ABC ltd. than project Y.

Project X has a standard deviation of 1,543,372.93 whereas project Y has a standard deviation of 4,582,575.70. Project Y thus has a higher standard deviation than project X. This means that project Y has a much higher market volatility than project X and thus a a higher risk to invest. Project X does not stray as far from its expected value, relative to Y, and thus not a great risk. However, project Y has a higher deviation from its expected return in the positive direction and may have a greater payout potential for ABC ltd.